OpenStudy (anonymous):

Solve the equation using the Zero-Product Property. (x-2)(x+7)=0 A. 2, -7 B. 2, 7 C. –1,1 D. –2, -7

OpenStudy (anonymous):

(x-2)(x+7)=0 if either (x-2)=0 or (x+7)=0 since anything multiplied by zero is zero so just solve x-2=0 and x+7=0

OpenStudy (anonymous):

u know i need to make u as my friend because u break it down lol

OpenStudy (anonymous):

so it B

OpenStudy (anonymous):

when you solve x+7 = 0 -7 -7 x=-7

OpenStudy (anonymous):

in other words a is ur answer

OpenStudy (anonymous):

already got it down ..lol